Two cars leave towns 750 kilometers apart at the same time and travel toward each other. One car's rate is 10 kilometers per hour less than the other's. If they meet in 5 hours, what is the rate of the slower car?
I have no idea how to even start this problem. Is there an equation or what do I do to solve this?

Answer by mananth(16946)

Car A x mph
Car B x+ 10
..
They are going away from each other.

So add up their speed.

combined speed = x+ 1 x + 10
( 2 x + 10 )
Time = 5 hours
Distance = 750 miles
Distance = speed * time
( 2 x 10 )* 5 = 750
10 x + 50 = 750
10 x = 750 + -50
10 x = 700
/ 10
x= 70 mph Car A
Car B 70 + 10 = 80 mph